7) General Mills common stock dividends have been growing at an annual rate of 7 percent per year over the past 10 years. Current dividends are $1.70 per share. What is the current value of a share of this stock to an investor who requires a 12 percent rate of return if the following conditions exist?

a. Dividends are expected to continue growing at the historic rate for the foreseeable future.

b. The dividend growth rate is expected to be 9 percent per year for the foreseeable future.

c. The dividend growth rate is expected to be 6.5 percent per year for the foreseeable future.
Answer #1

(a)-The current value of a share of this stock if the growth rate is 7.00%

Dividend per share in year 0 (D0) = $1.70 per share

Dividend Growth Rate (g) = 7.00% per year

Required Rate of Return (Ke) = 12.00%

The current value of a share of this stock = D0(1 + g) / (Ke – g)

= $1.70(1 + 0.07) / (0.12 – 0.07)

= $1.8190 / 0.05

= $36.38 per share

(b)-The current value of a share of this stock if the dividend growth rate is expected to be 9 percent per year for the foreseeable future

Dividend per share in year 0 (D0) = $1.70 per share

Dividend Growth Rate (g) = 9.00% per year

Required Rate of Return (Ke) = 12.00%

The current value of a share of this stock = D0(1 + g) / (Ke – g)

= $1.70(1 + 0.09) / (0.12 – 0.09)

= $1.8530 / 0.03

= $61.77 per share

(c)-The current value of a share of this stock if the dividend growth rate is expected to be 6.5 percent per year for the foreseeable future.

Dividend per share in year 0 (D0) = $1.70 per share

Dividend Growth Rate (g) = 6.50% per year

Required Rate of Return (Ke) = 12.00%

The current value of a share of this stock = D0(1 + g) / (Ke – g)

= $1.70(1 + 0.0650) / (0.12 – 0.0650)

= $1.8105 / 0.0550

= $32.92 per share